OpenStudy (katielong):

just a little indices query... *please wait whilst i write it as an equation* Thanks =)

OpenStudy (katielong):

express \[16^{3}\] as \[2^{?}\]

OpenStudy (katielong):

or express \[125^{4/3}\] as \[5^{?}\]

OpenStudy (anonymous):

first one is (2^4)=16 so replace 16 with 2^4-->(2^4)^3--->2^12=4096, which is the the same value as 16^3

OpenStudy (anonymous):

for the second one, i will only give you a hint: 125=5^3

OpenStudy (katielong):

sly hint... ahaha =)

OpenStudy (katielong):

how about expressing\[1/27\] as \[3^{?}\] ?

OpenStudy (apoorvk):

So, yes that was sly hint. Whaddya think? Try a guess! ;p

OpenStudy (anonymous):

125 \[125^{4/3} = (5^{3})^{4/3} = 5^{4}\]

OpenStudy (apoorvk):

hmm, do you know that, \[\large x^{-m} = \frac {1}{x^m} \]

OpenStudy (anonymous):

1/27 = 3^(-3)

OpenStudy (katielong):

thaanksss =)

OpenStudy (apoorvk):

Are you getting how we do this, the concept? Try to understand what LagrangeSon and Zeerak did up there^^.

OpenStudy (katielong):

yeahh thanks but i got an advanced one that im trying now... its hardd :/

OpenStudy (apoorvk):

Hmm, post that one too just in case you want to verify your answer :)

OpenStudy (katielong):

work out\[(1/64)^{-1/2}\]

OpenStudy (anonymous):

It is 8.

OpenStudy (katielong):

howww?? :S

OpenStudy (anonymous):

\[1/64^{-1/2} = 64^{1/2} = 8^{2*1/2} = 8\]

OpenStudy (katielong):

ahh i seee! thanks =)

OpenStudy (anonymous):

:)